a. How many moles of hydroxide ion are in a 100.0 L solution that has a pH of 10.00?

b. When the concentration of hydronium ion, [H3O+],decreases by a factor of 1000, what happens to the pH?

c. Which solution should have the greatest concentration of H3O+, and which should have the lowest pH? (One is a strong acid and one is a weak acid.)

0.10 M HNO2

0.10 M HNO3
Answer #1

i mole of OH- 10-4 x 100.0 pH = 10.00 poch = 14 - 10.0 POH= 4.0 -log [OH-] = 4.0 [OH-] = 104 m = 0.01 mole 6 pH = -log [143 o